If 360 g CO reacts with X g H2 then,

(i) Determine the grams of methanol that can be produced from these amounts of reactants.


1
Expert's answer
2021-01-04T03:46:21-0500

CO + 2H2 --> CH3OH

360g xg ??


Molar Mass of CH3OH = 12+3+16+1 = 32gmol-1


mole of CO = 360/28 = 12.86mol


mass of CH3OH = M.M * Mole

= 32gmol-1 * 12.86mol

= 411.52g